OpenStudy (anonymous):

On discovering that her family had a 70% risk of heart attack, Erin took a treadmill test to check her own potential of having a heart attack. The doctors told her that the reliability of the stress test is 67%. The test predicted that Erin will not have a heart attack. What is the probability after the test was taken that she will not have a heart attack? 0.099 0.201 0.231 0.469

OpenStudy (anonymous):

@Michele_Laino @ganeshie8

OpenStudy (michele_laino):

hint: we have to multiply the single probabilities. Namely the probability to have not heart attack is p_1=1-0.7=...? the probability to have not heart attack from the test, is p_2=1-0.67=...? so the requested probability, is: p_1*p_2=...?

OpenStudy (anonymous):

.3*.33?

OpenStudy (anonymous):

.099 as my answer? :)

OpenStudy (michele_laino):

sorry, please: p_2 =0.67

OpenStudy (anonymous):

.201 as my answer

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

Let's see

OpenStudy (anonymous):

it was wrong :(

ganeshie8 (ganeshie8):

try this (0.3*0.67)/(0.3*0.67+0.7*0.33)

OpenStudy (anonymous):

That gets .465

ganeshie8 (ganeshie8):

yes try last option

OpenStudy (anonymous):

Alright I will next try

ganeshie8 (ganeshie8):

btw that expression comes from conditional probability

ganeshie8 (ganeshie8):

P("no heart attack" | "test giving no heart attack") = P("no heart attack" AND "test giving no heart attack")/P("test giving no heart attack")

OpenStudy (michele_laino):

there is a p_2= 33% of probability to have no heart attack, and we have p_1=30% to have not heart attack. So try with: p_1*p_2=0.33*0.30=0.099

ganeshie8 (ganeshie8):

how can the probability of not having heart attack reduce from 30% to 9.9% even after the test showed no heart attack ?

ganeshie8 (ganeshie8):

actually we can eyeball the options without doing any work. the probability after the test has to be greater than 30% whatever that is. from your options only one is greater than 30% so..

OpenStudy (michele_laino):

I think that the test tell us that there is a 67% to have an heat attack, namely there is a 67% to confirm that 70% of probability to have an heat attack, so there is a 33% to have no heart attack from that test.

OpenStudy (anonymous):

It is not .469

ganeshie8 (ganeshie8):

oh

ganeshie8 (ganeshie8):

I might be wrong, but i don't really see how the probability of "not having heart attack" reduces even after the test predicts "no heart attack"

OpenStudy (anonymous):

Not sure

OpenStudy (anonymous):

Well it's not B, C, D so it must've been your first answer 0.099 :\ because I got it wrong thinking it was .231

OpenStudy (lyssakat):

ITS NOT A!!!!

OpenStudy (lyssakat):

NOT B